Weather Report for November 26, 2011

The storm of last night moved east very rapidly, leaving a photometric night in its wake. Seeing was generally poor during the first half of the night (around 2 arcsec) and improved to about 1.2-1.5 arcsec as morning approached.

About 1.5 hours were lost at the beginning of the night as a problem surfaced with SPOL. A fix was implemented, and all worked well for the rest of the night. Again, the problem was with the motion of the wave plate. Hopefully, the solution found tonight will be definitive.